OpenStudy (anonymous):

x% of y is z and (x + 10)% of y is equal to 2z, then y when expressed in terms of z is

OpenStudy (anonymous):

xy/10=z,(x+10)y/100=2z

OpenStudy (anonymous):

y=zx/100,yx+10y=200z

OpenStudy (anonymous):

zx/100+10zx/100=200z

OpenStudy (anonymous):

100z+10y^2-200zy

OpenStudy (dumbcow):

so we have: \[z = \frac{xy}{100} \rightarrow y = \frac{100z}{x}\] \[2z = \frac{y(x+10)}{100} = \frac{2xy}{100}\] From 2nd equation you can find that x=10 thus \[y = \frac{100z}{10} = 10z\]

OpenStudy (anonymous):

2xy/100 how?

OpenStudy (anonymous):

yx+10y/100

OpenStudy (dumbcow):

i substituted previous value for "z" and multiplied by 2

OpenStudy (anonymous):

A one line Mathematica solution is attached.

OpenStudy (anonymous):

oh ok

OpenStudy (dumbcow):

ahh sorry i realized i didn't make that clear \[2z = 2(\frac{xy}{100}) = \frac{y(x+10)}{100}\]

OpenStudy (anonymous):

k no problem

OpenStudy (anonymous):

thanks @robtobey